after a price increase of 10%, a television set costs $286 . what was the price before the increase.



Answer by Theo(13342)

after a 10% increase in the price, the price of the television set is 286.

let x equal the price of the television set before the increase.

you get x + .10 * x = 286.

factor out the x to get x * (1 + .10) = 286.

simplify to get x * 1.10 = 286

divide both sides of this equation by 1.10 to get x = 286 / 1.10.

solve for x to get x = 260.

the price of the television set was 260 before the increase of 10%.

10% of 260 = 26.

this is added to the price of the television set to get the new price of 260 + 26 = 286.